Suppose XYZ Company’s stock is currently selling for $30, has just paid a dividend of $3.00 and dividends are expected to grow at a constant rate of 4%. Also, assume that XYZ has bonds that are currently selling for $975.50 (par value is $1,000) with a coupon of 6% that is paid annually and a maturity of 15 years. If the tax rate is 21% and XYZ uses a capital structure of 30% debt and 70% equity, calculate the firm’s weighted average cost of capital.
Answer #1

The cost of equity capital is:

Re = D1/po + g

= 3*1.04/ 30 + 0.04

= 14.4%

the cost of debt is :

FV = $1000

PV = ($975.5)

PMT = $60

N = 15 YEARS

I/Y = 6.25

POST TAX COST OF DEBT IS = 6.25% * (1 - 0.21)

= 4.9375%

So, the WACC is :

weight of equity * cost of equity +  weight of debt* after tax cost of debt

=0.7* (0.144) + 0.3* ( 0.0494)

= 0.1008 + 0.0148

= 11.56%

So, the weighted average cost of capital is 11.56%.